n mole each of H2O, O2 and H2 are mixed at a suitable high temperature to attain the equilibrium

2H2O <--->2H2 +o2.

If Y mole of H2O are dissociated and the total pressure maintained is P, calculate hte Kp.

please explain every step.. thank you

Grade:12

1 Answers

Aman Bansal
10 years ago

Dear Student.

12.123 has 5 significant figures and 3 decimal places. 12.0123 has 6 significant and 4 decimal places. It is incorrect to take numbers apart and refer to individual portions as having their own significant figures. Here is a good example: The log of a product is equal to the sum of the logs of each multiplier: log(2.73 ×10^–5) = log(2.73) + log(10^–5) log(2.73) = 0.436: the answer has three significant figures, reflecting the possible error in the last digit of 2.73. log(10^–5) = –5.000000000...: the answer has an infinite number of significant digits because 10^–5 is an exact number and has no error. Then, log(2.73×10^–5) = log(2.73) + log(10^–5) = 0.436 + (–5.00000000...) requires that we use the rules for significant digits for sums, i.e., we can only add to the same decimal place as the value with the least significance. In this case, that is the third decimal place from 0.436, so 0.436 + (–5.000) = –4.564 or log(2.73 ×10^–5) = –4.564

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. 

http://www.askiitians.com/packages/packages.aspx

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert